IF IN THE SAMPLE OF HARD WATER PPM CONCENTRATION OF CALCIUM IS 400 PPM THE % OF CA IN THIS SAMPLE(WT/WT)

Grade:11

1 Answers

Aarti Gupta
9 years ago
As 1% = 1/100
and 1ppm = 1/1000000
so,
1 ppm = 0.0001%
now for 400 ppm of Ca = 0.0400%
To express in wt/wt
Mass % of Ca = Mass of Ca in the solution/ Total mass of the solution * 100
Mass % of Ca = 0.04/99.96 * 100
= 0.0004 * 100 = 0.04
Thanks & Regards
Aarti Gupta
askiitians Faculty